OpenStudy (anonymous):

Find the diagonal of a cube if its side equals 5. #SQR#

OpenStudy (tootsi123):

what are ur answer chioces

OpenStudy (tootsi123):

@JOEBARNES

OpenStudy (anonymous):

\[( ? ) \sqrt{?}\]

OpenStudy (tootsi123):



OpenStudy (tootsi123):

i'm working on this same thing so i know how it can be

OpenStudy (anonymous):

could it be \[2\sqrt{5}\]

OpenStudy (tootsi123):

I think that is right

OpenStudy (anonymous):

no the answer could be \[(9)\sqrt{8}\]